在一场有大量选手参加的 10 公里赛跑中,完赛时间服从正态分布。40 分钟后,有 10% 的选手完成了比赛。1 小时后,仍有 35% 的选手尚未完赛。最先完赛的前 20% 的选手将获得奖牌。

  1. (i)

    证明选手的跑步时间的均值为 55.4 分钟,标准差为 12.0 分钟(均精确到小数点后 1 位)。

  2. (ii)

    求选手为了能获得奖牌,完赛最多可以花多少时间。

  3. (iii)

    随机抽取 12 名选手。

    求超过 4 名选手获得奖牌的概率。

  4. (iv)

    已知这些选手均未获得奖牌,求最慢的选手在 1 小时内完赛的概率。

先不看答案,让 AI 诊断你卡在哪里?

它不会直接告诉你答案,而是问你卡在哪一步看看 AI 是怎么陪你一步步弄懂这道题的

完整解法

9(i)

1. Set up the probability statements and obtain standard normal equations using the inverse normal function.

Let TT be the running time in minutes. T∼N(μ,σ2)T \sim \mathrm{N}(\mu, \sigma^2).

Given 10% complete the race under 40 minutes:
P(T⩽40)=0.10  ⟹  40−μσ≈−1.28155\mathrm{P}(T \leqslant 40) = 0.10 \implies \frac{40-\mu}{\sigma} \approx -1.28155

Given 35% take more than 60 minutes:
P(T>60)=0.35  ⟹  P(T⩽60)=0.65  ⟹  60−μσ≈0.38532\mathrm{P}(T > 60) = 0.35 \implies \mathrm{P}(T \leqslant 60) = 0.65 \implies \frac{60-\mu}{\sigma} \approx 0.38532

2. Form simultaneous linear equations and solve for $\sigma$.

40−μ=−1.28155σ— (1)40 - \mu = -1.28155\sigma \quad \text{--- (1)}
60−μ=0.38532σ— (2)60 - \mu = 0.38532\sigma \quad \text{--- (2)}

Subtracting equation (1) from (2):
20=(0.38532−(−1.28155))σ20 = (0.38532 - (-1.28155))\sigma
20=1.66687σ20 = 1.66687\sigma
σ=201.66687≈11.9985\sigma = \frac{20}{1.66687} \approx 11.9985

3. Solve for $\mu$ and state the answers correct to 1 decimal place.

Substitute σ\sigma into equation (2):
60−μ=0.38532(11.9985)60 - \mu = 0.38532(11.9985)
60−μ≈4.623260 - \mu \approx 4.6232
μ=60−4.6232≈55.3768\mu = 60 - 4.6232 \approx 55.3768

Rounding to 1 decimal place:
μ≈55.4 minutes\mu \approx 55.4 \text{ minutes}
σ≈12.0 minutes\sigma \approx 12.0 \text{ minutes} (Shown)

答案

μ≈55.4\mu \approx 55.4 minutes, σ≈12.0\sigma \approx 12.0 minutes (Shown)

9(ii)

1. Set up the probability equation for the top 20% of runners who finish first.

Let TT be the time taken by a runner to finish the race.
The first 20% of runners are those with the shortest finishing times. We need to find the maximum time tt such that:
P(T<t)=0.2\mathrm{P}(T < t) = 0.2

2. Solve for the maximum time $t$ using the inverse normal distribution.

Using the unrounded parameters from part (i), μ≈55.377\mu \approx 55.377 and σ≈11.999\sigma \approx 11.999, we use the inverse normal function on the graphing calculator:
t≈45.279t \approx 45.279
t≈45.3 minutes (to 3 s.f.)t \approx 45.3 \text{ minutes (to 3 s.f.)}

答案

45.345.3 minutes

9(iii)

1. Define the distribution for the number of runners receiving a medal

Let XX be the number of runners in the sample of 12 who receive a medal.
Since the first 20% of runners receive a medal, the probability of any randomly selected runner receiving a medal is p=0.2p = 0.2.
X∼B(12,0.2)X \sim \mathrm{B}(12, 0.2)

2. Calculate the probability that more than four runners receive a medal

P(X>4)=1−P(X≤4)\mathrm{P}(X > 4) = 1 - \mathrm{P}(X \le 4)
Using a graphing calculator:
P(X>4)=1−0.92744\mathrm{P}(X > 4) = 1 - 0.92744
P(X>4)=0.072556≈0.0726 (to 3 s.f.)\mathrm{P}(X > 4) = 0.072556 \approx 0.0726 \text{ (to 3 s.f.)}

答案

0.07260.0726

9(iv)

1. Translate the scenario into conditional probability events for the sample of 12 runners.

Let AA be the event that none of the 12 runners receives a medal.
Let BB be the event that the slowest of the 12 runners completes the race in under 60 minutes.

For the slowest runner to take under 60 minutes, all 12 runners must take under 60 minutes.
We require P(B∣A)=P(A∩B)P(A)\mathrm{P}(B \mid A) = \frac{\mathrm{P}(A \cap B)}{\mathrm{P}(A)}.

2. Determine the required probabilities for a single runner using the exact percentages given in the problem stem.

Let TT be the completion time of a single runner, and tmt_m be the maximum time to receive a medal.
From the given information:
P(T<tm)=0.20\mathrm{P}(T < t_m) = 0.20 (the first 20% receive medals)
P(T>60)=0.35  ⟹  P(T≤60)=0.65\mathrm{P}(T > 60) = 0.35 \implies \mathrm{P}(T \le 60) = 0.65

For a single runner to not receive a medal, T>tmT > t_m:
P(T>tm)=1−0.20=0.80\mathrm{P}(T > t_m) = 1 - 0.20 = 0.80

For a single runner to not receive a medal AND complete in under 60 minutes, tm≤T<60t_m \le T < 60:
P(tm≤T<60)=P(T<60)−P(T≤tm)=0.65−0.20=0.45\mathrm{P}(t_m \le T < 60) = \mathrm{P}(T < 60) - \mathrm{P}(T \le t_m) = 0.65 - 0.20 = 0.45

3. Calculate the joint and marginal probabilities for the sample of 12 independent runners.

Since the 12 runners are randomly selected and independent:
P(A)=[P(T>tm)]12=0.8012\mathrm{P}(A) = [\mathrm{P}(T > t_m)]^{12} = 0.80^{12}
P(A∩B)=[P(tm≤T<60)]12=0.4512\mathrm{P}(A \cap B) = [\mathrm{P}(t_m \le T < 60)]^{12} = 0.45^{12}

4. Evaluate the final conditional probability.

P(B∣A)=0.45120.8012=(0.450.80)12=(0.5625)12≈0.000993 (to 3 s.f.)\begin{aligned} \mathrm{P}(B \mid A) &= \frac{0.45^{12}}{0.80^{12}} \\ &= \left(\frac{0.45}{0.80}\right)^{12} \\ &= (0.5625)^{12} \\ &\approx 0.000993 \text{ (to 3 s.f.)} \end{aligned}

(Note: Using the rounded normal distribution parameters μ=55.4\mu = 55.4 and σ=12.0\sigma = 12.0 from part (i) yields a slightly inaccurate 0.0009740.000974, which typically loses an accuracy mark. Relying on exact given percentages is the intended method.)

答案

0.0009930.000993

AI 分析,可能出错

用 AI 批改,圈出写错的那一步,并自动将你的薄弱点存入错题本,形成你的错因库。

同一知识点的其他题

  1. 1永初农场(Eng Choo Farm)的业主声称,其农场生产的鸡蛋平均每只含有 μ0\mu_0 毫克的胆固醇。 健康促进局从该农场随机抽取了 16 只鸡蛋,并测量了每只鸡蛋的胆固醇含量 xx 毫克。结果汇总如下: ∑x=2848\sum x = 2848…
  2. 2在本题中,你应该写出你所使用的任何分布的参数。 Crispy Cream 甜甜圈店销售 2 种甜甜圈:环形甜甜圈(Ring Donuts)和夹心甜甜圈(Filled Donuts)。环形甜甜圈和夹心甜甜圈的质量(以克为单位)分别服从正态分布…
  3. 3XX 是一个均值为 1、方差为 1 的正态随机变量,并且 YY 也是一个均值为 μ\mu、方差为 2 的正态随机变量。 已知 P(0≤X+Y<3)>0.44P(0 \le X + Y < 3) > 0.44,求 μ\mu 的所有可能值的范围。 若改为已知 μ=10\mu = 10…
  4. 4捕鱼公司 A 有两种类型的渔船:“标准型”渔船和“大型”渔船。在一次典型的捕鱼航行中,各渔船捕获的鱼量均呈正态分布,其均值和标准差如表所示。 捕鱼公司 B 与捕鱼公司 A 在同一海域作业。 然而,公司 B 拥有公司 A 两种渔船的改进版本:…
  5. 5在本题中,你应该清楚地陈述所使用的所有分布以及相应参数的值。 一位玩具设计师设计了用于平衡类益智玩具的加重木块。木块有两种类型:圆柱体木块和长方体木块。每个圆柱体木块和长方体木块的质量(以克为单位)均服从正态分布,参数如下: | | 均值…